OpenStudy (brooke..help00):

PLEASE SOMEONE HELP ME!!!!????????? WILL FAN ANYONE WHO ATTEMPTS AND MEDAL WHO HELPS ME THE QUICKEST!!!! COMMENT IF YOU CAN!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (will.h):

And the question is

OpenStudy (brooke..help00):

The table shows the solution to the equation |2x − 3| − 1 = 2: Step 1: |2x − 3| = 2 + 1 Step2: |2x − 3| = 3 Step 3: 2x − 3 = 3 or 2x + 3 = 3 Step 4: 2x = 6 or 2x = 0 Step 5: x = 3 or x = Which is the first incorrect step? Step 1 Step 3 Step 5 Solution is correct

OpenStudy (brooke..help00):

@Will.H

OpenStudy (will.h):

I'm wondering the original equation is |2x - 3| = 2 or 3?

OpenStudy (brooke..help00):

2

OpenStudy (1davey29):

Step 3. abs(2x-3)=3 would become 2x-3=3 or 2x-3=-3 because abs(x)=y is simplified to x=+y or -y

OpenStudy (will.h):

Step 1 on the right side why were there 2+ 1 if the side was only equal to 2

OpenStudy (will.h):

I don't find it make sense to add 1

OpenStudy (1davey29):

You would add the 1 to cancel the -1 when solving for x

OpenStudy (will.h):

Oh I didn't see that -1 actually